A car travels on a straight road with constant speed and its wheel rotates 450 rotations

in 12 s. If the diameter of the wheel of the car is 50 cm, calculate the

(a) angular speed of the wheel in SI unit.


(b) speed of the car. 


1
Expert's answer
2021-10-14T18:38:45-0400

a)

"\omega=2\pi\nu=2\pi\frac nt=235.5~\frac{rad}s,"

b)

"v=\omega r=\frac{\omega d}2=58.9~\frac ms."